#2: A Saturn-V Rocket has a total mass of 3.00 x 106 kg when fully fueled. The first stage of rocket propulsion burns 2.27 x 100 kg of fuel in 150 s, providing a thrust of 3.402 x 10 N, as the rocket is lifted against Earths gravity. a) Find the rate at which fuel is burned during this first-stage propulsion, assuming a steady rate of fuel consumption. b) Find the exhaust speed of the fuel as it is ejected from the rocket. c) Find the acceleration of the rocket when the rocket first starts to burn. d) Find the acceleration of the rocket immediately before the fuel is exhausted. Note: you must include the gravitational force to solve parts cand d
Answer #1

a) rate of burning=( 2.27 x 10^6) kg/ 150 sec = 1.5133 x 10^ 5 kg/sec

b) V+Ve ( exhaust speed of fuel with respect to rocket) = 3.402 x 10^7 ( 1/ 1.5133 x 10^ 5) = 2.24806 x 10^2 m/sec

c) a= ( 2.24806 x 10^2 )( 1.5133 x 10^ 5 kg/sec)/ 3 x 10^6 -9.8

a=1.54 m/s^2

d)let's calculate the final velocity ( the time at which the entire fuel is burned)'

v= 2.24806 x 10^2 ln (3 x1 0^6/ 0,.73 x 10^6) = 2.24806 x 10^2 x 1.413 = 3.178 x 10^2 m/s

a= 3.178 x 10^2/ 150= 2.11 m/s^2